OpenStudy (anonymous):

A wooden garden tray with a base area of xy square inches is filled with soil up to a certain height. The tray contains (2xy + xy2 + x2y2) cubic inches of soil. The height of the tray is represented by the expression (2xy+xy2+x2y2)/(xy) Which expression below is the correct simplified form of the height of the tray? 2x3y4 2x3y3 2 + xy2 + xy 2 + y + xy idk this one, I need help

OpenStudy (anonymous):

A news agent conducted a survey among travel magazine subscribers of a town and found that there was circulation of only three travel magazines. He made the Venn diagram below to show the number of subscribers to each of the three magazines: Leisure, Outbound, and Vacations. image link: https://www.connexus.com/content/media/369615-772011-32443-PM-1657781887.png How many subscribers belong to the set V U (~O ∩ L)? 749 821 870 954

OpenStudy (anonymous):

there's also this question...

OpenStudy (anonymous):

these are my last two questions... they are the only two I didnt have an answer for...

OpenStudy (anonymous):

can anyone help me out?

OpenStudy (anonymous):

when i did the first one, I got xy+y+2?

OpenStudy (anonymous):

would that mean its d?

OpenStudy (anonymous):

D is my final answer for the first one, is that correct?

OpenStudy (anonymous):

answer for the first one is D you got it correct

OpenStudy (anonymous):

koay, cool now my last question of the day... dun dun dun.... lol

OpenStudy (anonymous):

okay*

OpenStudy (anonymous):

i m doing ur venn diagram question

OpenStudy (anonymous):

okay, thanks so much :)

OpenStudy (anonymous):

I swear I must be crazy, ive been working on the same pre examination test for the past 6 hours... ugh math is so tiring... lol im glad this is my last question, but then I have to go and take the actual test.... can't wait :/ lol

OpenStudy (anonymous):

ok actually maths needs alot of practice

OpenStudy (anonymous):

ohhh, I know! you have to keep at it or else you will forget EVERYTHING! its happened to me and I had to relearn Algebra 1 twice... lol

OpenStudy (mertsj):

Which one did you message me about?

OpenStudy (anonymous):

the Venn diagram one, with the link attached...

OpenStudy (anonymous):

i m good at maths as i m a graduate student i do alot of practice i can solve question while watching T.V

OpenStudy (anonymous):

what? damn! how can you manage to do that. I have to isolate myself in my room in order to be able to focus... lol

OpenStudy (anonymous):

that's just amazing! :)

OpenStudy (anonymous):

r u taking complement with O???

OpenStudy (anonymous):

idk, im not sure what the squiggle means... sorry :(

OpenStudy (anonymous):

I think it means similar? or something idk??

OpenStudy (anonymous):

like in proportion....

OpenStudy (anonymous):

can just repost the actually equation once again cauz it making trouble in answer

OpenStudy (anonymous):

no problem, here it is: A news agent conducted a survey among travel magazine subscribers of a town and found that there was circulation of only three travel magazines. He made the Venn diagram below to show the number of subscribers to each of the three magazines: Leisure, Outbound, and Vacations. image link: https://www.connexus.com/content/media/369615-772011-32443-PM-1657781887.png How many subscribers belong to the set V U (~O ∩ L)? 749 821 870 954

OpenStudy (anonymous):

This is what I have so far: V=387+67+72=526 O∩L = 84+49=133 526+133=659 that was my guess, buy I know im doing something wrong, I dont have a reference sheet, so Im unsure on how to solve it... :/

OpenStudy (anonymous):

yeah i m getting the same answer... we have to take a complement of O

OpenStudy (anonymous):

@Mertsj, what do you think, im stumped...

OpenStudy (anonymous):

idk, what should I put?

OpenStudy (anonymous):

659 doesn't match with any of the options... :(

OpenStudy (anonymous):

trust me its really very close to the answer

OpenStudy (anonymous):

would it be A? that's the closest one

OpenStudy (anonymous):

isnt it funny how the last question always seems the hardest... lol

OpenStudy (anonymous):

but yeah, @Mertsj, do you have any idea? did we do the problem correctly or are we missing something?

OpenStudy (anonymous):

????

OpenStudy (anonymous):

@cshalvey its the one with the link

OpenStudy (anonymous):

I think i mite have it, (~O ∩ L) is everything in L but outside of O. That would be the two regions of L with values 295 and 67. When we apply the union operation, we add any regions of V that have not already been included. That would be the regions of 72 and 387 and 49 . Adding, we get 295 + 67 + 72 + 387 + 49. = 870. Does that sound right?

OpenStudy (anonymous):

yeah its making logic

OpenStudy (anonymous):

Yeah, I think you've got it there - I'm coming up with the same thing...

OpenStudy (anonymous):

yap its 870

OpenStudy (anonymous):

yay! :) okay thanks you guys I have to go submit it now, ill be back on in a little bit :)

OpenStudy (anonymous):

Good work @annas and @Renee99 (think you ended up solving it all by yourself) ;). Way to go both of you - great to see this kind of interaction

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

anytime and thanks cs really appreciate your help

OpenStudy (mertsj):

I got 870

OpenStudy (anonymous):

thanks again to all of you :)